OpenStudy (lindamathmath):

Raina the trainer has two solo workout plans that she offers her clients: Plan A and Plan B. Each client does either one or the other (not both). On Monday there were 9 clients who did Plan A and 7 who did Plan B. On Tuesday there were 3 clients who did Plan A and 5 who did Plan B. Raina trained her Monday clients for a total of 12 hours and her Tuesday clients for a total of 6 hours. How long does each of the workout plans last?

alones (alones):

oh this looks so confusing /.<

OpenStudy (lindamathmath):

Raina the trainer has two solo workout plans that she offers her clients: Plan A and Plan B. Each client does either one or the other (not both). On Monday there were 9 clients who did Plan A and 7 who did Plan B. On Tuesday there were 3 clients who did Plan A and 5 who did Plan B. Raina trained her Monday clients for a total of 12 hours and her Tuesday clients for a total of 6 hours. How long does each of the workout plans last?

OpenStudy (lindamathmath):

whoever helps me gets a cookie

jimthompson5910 (jim_thompson5910):

`On Monday there were 9 clients who did Plan A and 7 who did Plan B` 9+7 = 16 clients total for Monday `Raina trained her Monday clients for a total of 12 hours` so, time per client = (total time)/(number of clients) time per client = 12/16 time per client = 3/4 hours = 45 minutes basically, each client took 45 minutes to train This is all for Monday. Do the same steps for Tuesday.

OpenStudy (lindamathmath):

thanks boi